40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 17 03:30:13 2012, Den, Ayaka, Update, Alignment, red in arms 
    Reply  Sat Nov 17 18:02:14 2012, Den, Update, Alignment, red in arms 
       Reply  Sat Nov 17 22:50:17 2012, Koji, Update, Alignment, red in arms 
       Reply  Mon Nov 19 15:12:52 2012, Jenne, Update, Alignment, Yarm locked IR 
Message ID: 7723     Entry time: Mon Nov 19 15:12:52 2012     In reply to: 7721
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: Yarm locked IR 

Quote:

Quote:

POY11 does not go out of the vacuum

 It does but slighty low and does not get on mirrors. We need to change optic mounts to adjust the height. Red is flashing in yarm at 00 and 10 modes. TRY is ~0.4-0.5.

I've adjusted BS angle, camera and TRX PD at ETMX table so I can see red flashing at 03 mode while green is locked to 00 and its transmission is maximized. I thought that by adjusting BS angle, I will be able to align red to 00 not disturbing green, but this was not the case. Maximum TRX I could get was 0.1. I've adjusted POX to get into PD and I can see PDH signal though I can't lock as cavity is still misaligned for red.

 [Ayaka, Jenne]

We put the POY beam onto the POY PD.  The Yarm is currently locked on IR with ~0.65 transmission.

 

ELOG V3.1.3-